We have to measure the distance from the diagonal of our computer screen and to the nearest tenth of centimeter or sixteenth of an inch. Measure the height of the screen along the vertical as well. I came up the measurements of 19 for the diagonal and 12 for the height and 14 for the horizontal. Could you please help me solve? I measured and got 14 for my height and my measurements are off? Please help
Answer by stanbon(75887)

We have to measure the distance from the diagonal of our computer screen and to the nearest tenth of centimeter or sixteenth of an inch. Measure the height of the screen along the vertical as well. I came up the measurements of 19 for the diagonal and 12 for the height and 14 for the horizontal. Could you please help me solve? I measured and got 14 for my height and my measurements are off? Please help
--------------
If you measured the diagonal as 19" and the vertical as "12"
Pythagoras can figure out the horizontal, if that is what you
want to know:
19^2 = 12^2 + (horizontal)^2
horizontal = sqrt(19^2 - 12^2)
horizontal = sqrt(217)
horizontal = 14.7309"
==================================
Cheers,
Stan H.